# The impact of $\gamma N$ and $\gamma^{*}N$ interactions on our understanding of nucleon excitations

### v1

@@ 4.5   Functional methods
Progress in calculating the baryon spectrum has also been made with functional methods. Here the central quantities of interest are QCD’s $n$-point correlation functions in Fig. 4.2. As mentioned in Sec. 4.1, they can be obtained through functional derivatives of the QCD partition function (4.5). Taking the quantum expectation values of the classical (Klein-Gordon, Maxwell, Dirac) equations of motion yields the Dyson-Schwinger equations (DSEs), which relate different $n$-point functions with each other [724, 658, 725, 726, 727, 692, 728, 729]. Similar systems of equations can be derived using the functional renormalization group (FRG) [730, 731, 732]. Therefore, in principle functional methods are ab-initio methods just like lattice QFT: while on the lattice one calculates the correlation functions directly from the path integral, the strategy with functional methods is to calculate them from each other using equations derived from the path integral.
@@
As examples, the left plot in Fig. 4.12 shows the DSEs for the quark and gluon propagator and the three-gluon vertex. Like usual Feynman diagrams, each loop represents a four-momentum integration. Because the quantities on the left appear again on the right inside the integrals, these are integral equations which can be solved recursively. In this process, the classical propagators and vertices from the Lagrangian get ‘dressed’ and acquire structure. For example, to solve the quark DSE one needs to know the gluon propagator and quark-gluon vertex. These satisfy their own DSEs, which depend again on the quark propagator but also on higher $n$-point functions. This yields an infinite tower of equations, which for practical calculations must be truncated. The truncations employed in the literature so far cover a broad landscape ranging from simple models to ab-initio calculations without any model input. For example, one could absorb all information from the gluon sector into an effective interaction and solve the quark DSE as a standalone equation. Or one could also solve the gluon DSE and set all three- and four-point functions to their classical values, or calculate all two- and three-point functions while neglecting four-point functions, etc. Therefore, systematically enlarging these equations should improve the precision, as is already visible in the existing DSEs and FRG calculations [733, 728, 734, 735, 736, 737]; e.g., recent glueball calculations in Yang-Mills theory (i.e., QCD without quarks) are in good agreement with lattice results [738, 739, 740]. Correlation functions with quark and gluon legs are gauge-dependent, so the strategy is to choose the best-suited gauge for calculating gauge-invariant observables – in analogy to choosing spherical coordinates to compute the energy levels of the hydrogen atom in quantum mechanics. Most calculations so far have employed the Landau gauge as it preserves manifest Lorentz invariance and is a fixed point under the renormalization group.
@@
Figure 4.12: (color online) Left: Dyson-Schwinger equations for the quark and gluon propagator and three-gluon vertex. The solid and dashed lines represent quarks and ghosts, the springs are gluons. Two-loop diagrams in the three-gluon DSE are omitted. Right: Quark mass function for different quark flavors obtained from the quark DSE [692].
@@
One advantage of functional methods is that they facilitate a ‘hands-on’ understanding of QCD’s nonperturbative phenomena. For example, a key feature of QCD is dynamical mass generation: The sum of the light quark masses only amounts to 1% of the proton mass, so the rest must somehow be generated in QCD. This is a direct outcome of the quark DSE: The calculated quark masses run with the momentum scale (right panel in Fig. 4.12), so that for low momenta the current-quark masses (which are inputs to QCD) turn into large constituent-quark masses around 350 MeV. For heavier flavors the effect is eventually washed out. This is a nonperturbative effect due to the dynamical breaking of chiral symmetry, which cannot be generated at any order in perturbation theory (see e.g. [741] for a pedagogical discussion). The resulting mass function at low momenta is proportional to the quark condensate and sets the scale for the pion decay constant as well as the masses of the nucleon and other hadrons. The dynamical generation of a quark mass function is also seen in lattice QCD [742, 743], where it arises from center vortices [744, 745]. Note that the influence of such topological gluon field configurations on the $n$-point correlation functions is not directly accessible through DSEs, which are derived from the partition function where the gluon fields are already integrated over – as mentioned above, in contrast to lattice calculations, DSEs determine the correlation functions from each other. Another robust prediction from both functional methods and lattice calculations is the disappearance of the massless pole in the gluon propagator, i.e., the gluon develops a mass gap, see [658, 746, 747, 748, 749, 750, 733, 728, 751, 735, 752] and references therein.
@@
Hadrons appear as poles in QCD’s correlation functions, like the $qqq$ correlation function in the second row of Fig. 4.3. An efficient way to extract the mass of a hadron, like e.g. a nucleon resonance, is again the Bethe-Salpeter equation (BSE) discussed in Sec. 4.2, but now formulated in terms of quarks and gluons. One writes down a scattering equation $\mathbf{T}=\mathbf{K}+\mathbf{K}\,\mathbf{G}_{0}\,\mathbf{T}$, where $\mathbf{T}$ is now the connected and amputated part of the $qqq$ correlation function (i.e., without external quark propagators). $\mathbf{G}_{0}$ is the product of three dressed quark propagators which depend on the quark mass function in Fig. 4.12, so the quarks are also no longer onshell particles with a simple mass pole and a three-dimensional reduction is no longer possible. Because $\mathbf{T}$ is amputated, its residue at a given hadron pole is the hadron’s Bethe-Salpeter amplitude $\mathbf{\Gamma}$ defined through $\Psi=\mathbf{G}_{0}\,\mathbf{\Gamma}$, where $\Psi$ is the Bethe-Salpeter wave function (BSWF) from Eq. (4.8). Comparing the residues on both sides of the scattering equation then yields the homogeneous BSE at the pole,
@@
The wave function has three Dirac indices for the quarks and one for the baryon, and depending on the spin $J$ there may also be further Lorentz indices. It depends on three independent momenta, e.g. two relative Jacobi momenta $p$ and $q$ and the total momentum $P$. The wave function can be expanded in a Lorentz-covariant tensor basis, which yields 64 linearly independent tensors $\tau_{i}(p,q,P)$ for $J=1/2$ states and 128 for $J=3/2$ states [753, 754]. From three independent momenta one can form six Lorentz invariants, one of them being onshell ($P^{2}=m^{2}$), so the dressing functions $f_{i}$ depend on five momentum variables.
@@
What are the new effects compared to the quark-model wave functions discussed in Sec. 3.2? One can arrange the $\tau_{i}$ in Eq. (4.35) into eigenstates of the three-quark spin $S$ and orbital angular momentum $L$ in the baryon’s rest frame, which results in tensors with $L=0$ ($s$ waves), $L=1$ ($p$ waves), $L=2$ ($d$ waves), and so on [692]. These components mix together in the BSWF to produce a baryon with definite $J^{P}$. For example, the nucleon flavor wave function is a doublet $\mathcal{D}_{f}$, so one can cast Eq. (4.35) in permutation-group doublets $\mathcal{D}_{i}$ to obtain a fully symmetric wave function $\Psi=\sum_{i=1}^{64}\mathcal{D}_{i}\cdot\mathcal{D}_{f}$ (modulo color). One of these components is $\psi_{1}=(\mathcal{S}_{o}\,\mathcal{D}_{s})\cdot\mathcal{D}_{f}$ from Fig. 3.1, which we earlier identified with ‘the’ nucleon wave function. However, there are other $L=0$ contributions which are similarly important, and relativistic $p$ waves with $L=1$ contribute another substantial fraction of about 33% to the nucleon’s normalization [755], as we will discuss below in connection with Fig. 4.14. Relativity is therefore important for light baryons, and relativistic effects even persist up to bottom baryons [756].
@@
The homogeneous BSE (4.34) is equally applicable for bound states and resonances, just like Eq. (4.8) holds in both cases. To extract a resonance mass, however, one must analytically continue the eigenvalues to unphysical Riemann sheets, which can be technically challenging [757, 758]. The situation is also more involved than with EFTs, where the particles in the loops are hadrons and so the multihadron cuts are automatic. In Fig. 4.13, the particles in the loops are quarks and gluons, so the BSE must simultaneously create hadron poles dynamically from the quark level but also the multihadron cuts, which create the Riemann sheet structure to allow for resonances, all while avoiding decay channels for quarks and gluons as required by confinement. Such a mechanism has been demonstrated in Ref. [759]: A four-quark BSE implementing gluon exchanges dynamically creates intermediate meson poles, and these automatically also generate meson-meson cuts which turn a four-quark ($qq\bar{q}\bar{q}$) state into a resonance.
@@
Figure 4.14: (color online) Top: Light baryon spectrum from the three-quark and quark-diquark calculations for different $J^{P}$ quantum numbers [760, 761], compared to the RPP masses (real parts of pole positions). Bottom: Orbital angular momentum contributions to each baryon from the quark-diquark calculation. The bars sum up to 100%, and the light-colored boxes drawn around them are the quark-model values for $L$ from Fig. 3.2.
@@
What is the kernel $\mathbf{K}$? This is where the information from QCD’s $n$-point function enters. The leading term in the irreducible three-body kernel in Fig. 4.13, which is the three-gluon vertex connecting the three quarks, vanishes by the color traces. This hints at quark-quark correlations as the dominant forces in baryons. For the two-body kernels, most studies so far have used the rainbow-ladder truncation, where the BSE kernel turns into an effective gluon exchange [762, 763, 764]. This preserves chiral symmetry and its dynamical breaking pattern: The quark DSE dynamically generates quark masses like in Fig. 4.12, and the pseudoscalar mesons calculated from their BSEs are the massless Goldstone bosons in the chiral limit. In the light-meson sector, rainbow-ladder works well for pseudoscalar and vector mesons, whereas scalar and axialvector mesons are too strongly bound. Going beyond rainbow-ladder by implementing higher $n$-point functions is computationally demanding, but efforts have been made and improved the meson spectrum [765, 766, 692, 738].
@@
Concerning the baryon spectrum, one strategy has been to solve the three-body equation in rainbow-ladder without further approximations [753, 755, 754, 767, 756, 768, 769]. The other is to simplify the three-body equation to a quark-diquark BSE shown in the right panel of Fig. 4.13, where gluons no longer appear and the interaction becomes a ping-pong quark exchange, with quarks forming (non-pointlike) diquarks in the process [770, 771, 760, 692, 772, 773, 522, 774]. Fig. 4.14 shows the resulting light baryon spectrum [760], which is very similar in both cases. In particular, the nucleon mass comes out at $940$ MeV in both approaches without any tuning. Diquark clustering is a good way to think about baryons!
@@
An interesting observation in Fig. 4.14 is the correct level ordering between the Roper resonance $N(1440)\nicefrac{{1}}{{2}}^{+}$ and the nucleon’s parity partner $N(1535)\nicefrac{{1}}{{2}}^{-}$. This has not so much to do with the Roper resonance but rather with the $N(1535)$, which in contrast to the Roper is very sensitive to the higher-lying (‘ugly’) pseudoscalar and vector diquarks. In the rainbow-ladder calculation the $N(1535)$ mass comes out too low, along with many other states in the spectrum. If the strength of these ‘ugly’ diquarks is reduced, the $N(1535)$ mass moves up, overtakes the Roper resonance and ends up in the proximity of the mass of the $N(1650)$ (which is also not sensitive to the higher-lying diquarks). Corresponding plots can be found in Ref. [522]. In the three-body calculation, where diquarks do not appear explicitly, this means that the $N(1535)$ is very sensitive to effects beyond rainbow-ladder and, therefore, gluon exchange alone is not sufficient to describe the $N(1535)$ well. This is analogous to calculations in the meson sector, where the scalar and axialvector mesons receive large repulsive corrections beyond rainbow-ladder [765, 766]. Note also that the observed overall agreement with experiment in Fig. 4.14 is not in conflict with the importance of meson-baryon interactions, which are obviously necessary to generate a resonance mechanism. Unlike for higher multiquark states, such a mechanism is not automatic in the three-quark system and would enter through more advanced truncations with quark and gluon correlation functions that produce internal meson and baryon poles. The ratio $M/f_{\pi}$ (which is essentially what is shown in Fig. 4.14) would also remain unchanged if meson-baryon interactions affected the masses and $f_{\pi}$ proportionally. In any case, as expected from our discussion in Sec. 4.1, these results imply that all nucleon excitations below 2 GeV have a three-quark core.
@@
It is also interesting to study the internal composition of these baryon excitations [760, 761, 772, 775, 774]. The bars in the lower panel of Fig. 4.14 show the orbital angular momentum contributions to each baryon’s normalization. These come from the various tensors contributing to the wave functions discussed in connection with Eq. (4.35), i.e., each baryon is a mixture of different $L$ components in the rest frame. For comparison, the light-colored boxes show the quark-model values for $L$ in the harmonic oscillator assignment according to Fig. 3.2, where each baryon carries a specific value of $L$. For the majority of states one can see clear traces of the nonrelativistic quark model, e.g., the nucleon states with $J^{P}=1/2^{-}$ are dominated by $L=1$ and those with $J^{P}=3/2^{+}$ by $L=2$. But there are also clear deviations: The nucleon and $\Delta(1232)$ ground states are dominated by $L=0$ but have substantial (relativistic) $L=1$ contributions, and for several states the nonrelativistic assignment does not work at all.
@@
As will be discussed in Sec. 5, further information on the internal structure of nucleon resonances can be extracted from nucleon-to-resonance transition form factors. With functional methods, these have been studied in a variety of different approximations ranging from contact interactions [776, 777, 778, 779, 780], the quark-diquark approach either with model input [781, 782, 783, 784, 785] or calculated rainbow-ladder ingredients [755], or in the three-quark approach [786]. Like for the spectrum, the three-quark and quark-diquark calculations show good overall agreement, see the reviews [692, 522]. A typical common feature is missing strength at low $Q^{2}$, which in some cases can be clearly attributed to missing meson-cloud effects, e.g. for the nucleon elastic electromagnetic form factors: In that case, the meson effects are enhanced in the isovector combination but cancel in the isoscalar one, and the isoscalar anomalous magnetic moment from the three-body equation is in agreement with the experimental result $\kappa_{s}=-0.12$ [755]. In other cases, discrepancies that are sometimes attributed to the meson cloud can be explained by the relativistic $L=1$ components in the baryon wave functions, like for the $R_{EM}$ ratio in the $N\to\Delta\gamma$ transition [787]. For the Roper resonance, the transition form factor $F_{2}$ shows the zero crossing expected from a radial excitation, which points towards an underlying three-quark composition [781]. Especially important in the timelike region ($Q^{2}<0$) are the vector-meson poles, which are part of the dressed quark-photon vertex and automatically generated from the quark-gluon interactions if the vertex is calculated self-consistently [788]. More advanced truncations that are currently being explored also capture the $\rho\to\pi\pi$ resonance mechanism which turns the $\rho$ meson into a resonance [789, 790].
@@
Similar plots as Fig. 4.14 can be drawn for hyperons [791, 792] and heavy baryons [756, 768, 793]. Especially interesting are charmed baryons such as $\Sigma_{c}$ and $\Lambda_{c}$, which consist of two light quarks and one charm quark. Denoting up and down by $n$, it is often assumed that charmed baryons form ‘planetary’ systems made of a light $(nn)$ diquark that ‘orbits’ around the heavier charm quark. However, the Faddeev calculations do not support this picture [794, 793]. Rather, the $(nc)n$ configurations are dominant and contribute about $60\%$ to the norm of the state, while the $(nn)c$ configurations contribute only $40\%$. Therefore, a charmed baryon actually spends most of its time in a $(nc)n$ configuration.
@@
Figure 4.15: (color online) Left: Comparison of baryon octet and decuplet masses from different lattice calculations [795, 796, 797, 798, 799]. Plot from Ref. [799]. Right: Lattice finite-volume energy levels for $N^{*}$ and $\Delta^{*}$ excitations at a pion mass of 396 MeV [800].

### v4

@@ 4.5  Functional methods
Progress in calculating the baryon spectrum has also been made with functional methods. Here the central quantities of interest are QCD’s $n$-point correlation functions in Fig. 4.2. As mentioned in Sec. 4.1, they can be obtained through functional derivatives of the QCD partition function (4.5). Taking the quantum expectation values of the classical (Klein-Gordon, Maxwell, Dirac) equations of motion yields the Dyson-Schwinger equations (DSEs), which relate different $n$-point functions with each other Roberts:1994dr , Alkofer:2000wg , Fischer:2006ub , Bashir:2012fs , Cloet:2013jya , Eichmann:2016yit , Huber:2018ned , Fischer:2018sdj . Similar systems of equations can be derived using the functional renormalization group (FRG) Berges:2000ew , Pawlowski:2005xe , Dupuis:2020fhh . Therefore, in principle functional methods are ab-initio methods just like lattice QFT: while on the lattice one calculates the correlation functions directly from the path integral, the strategy with functional methods is to calculate them from each other using equations derived from the path integral.
@@
As examples, the left plot in Fig. 4.12 shows the DSEs for the quark and gluon propagator and the three-gluon vertex. Like usual Feynman diagrams, each loop represents a four-momentum integration. Because the quantities on the left appear again on the right inside the integrals, these are integral equations which can be solved recursively. In this process, the classical propagators and vertices from the Lagrangian get ‘dressed’ and acquire structure. For example, to solve the quark DSE one needs to know the gluon propagator and quark-gluon vertex. These satisfy their own DSEs, which depend again on the quark propagator but also on higher $n$-point functions. This yields an infinite tower of equations, which for practical calculations must be truncated. The truncations employed in the literature so far cover a broad landscape ranging from simple models to ab-initio calculations without any model input. For example, one could absorb all information from the gluon sector into an effective interaction and solve the quark DSE as a standalone equation. Or one could also solve the gluon DSE and set all three- and four-point functions to their classical values, or calculate all two- and three-point functions while neglecting four-point functions, etc. Therefore, systematically enlarging these equations should improve the precision, as is already visible in the existing DSEs and FRG calculations Cyrol:2017ewj , Huber:2018ned , Gao:2021wun , Eichmann:2021zuv , Fu:2022uow , Ihssen:2024miv ; e.g., recent glueball calculations in Yang-Mills theory (i.e., QCD without quarks) are in good agreement with lattice results Huber:2020ngt , Huber:2021yfy , Huber:2025kwy . Correlation functions with quark and gluon legs are gauge-dependent, so the strategy is to choose the best-suited gauge for calculating gauge-invariant observables – in analogy to choosing spherical coordinates to compute the energy levels of the hydrogen atom in quantum mechanics. Most calculations so far have employed the Landau gauge as it preserves manifest Lorentz invariance and is a fixed point under the renormalization group.
@@
Figure 4.12: (color online) Left: Dyson-Schwinger equations for the quark and gluon propagator and three-gluon vertex. The solid and dashed lines represent quarks and ghosts, the springs are gluons. Two-loop diagrams in the three-gluon DSE are omitted. Right: Quark mass function for different quark flavors obtained from the quark DSE Eichmann:2016yit .
@@
One advantage of functional methods is that they facilitate a ‘hands-on’ understanding of QCD’s nonperturbative phenomena. For example, a key feature of QCD is dynamical mass generation: The sum of the light quark masses only amounts to 1% of the proton mass, so the rest must somehow be generated in QCD. This is a direct outcome of the quark DSE: The calculated quark masses run with the momentum scale (right panel in Fig. 4.12), so that for low momenta the current-quark masses (which are inputs to QCD) turn into large constituent-quark masses around 350 MeV. For heavier flavors the effect is eventually washed out. This is a nonperturbative effect due to the dynamical breaking of chiral symmetry, which cannot be generated at any order in perturbation theory (see e.g. Eichmann:2025wgs for a pedagogical discussion). The resulting mass function at low momenta is proportional to the quark condensate and sets the scale for the pion decay constant as well as the masses of the nucleon and other hadrons. The dynamical generation of a quark mass function is also seen in lattice QCD Bowman:2005vx , Oliveira:2018lln , where it arises from center vortices Biddle:2023lod , Kamleh:2023gho . Note that the influence of such topological gluon field configurations on the $n$-point correlation functions is not directly accessible through DSEs, which are derived from the partition function where the gluon fields are already integrated over – as mentioned above, in contrast to lattice calculations, DSEs determine the correlation functions from each other. Another robust prediction from both functional methods and lattice calculations is the disappearance of the massless pole in the gluon propagator, i.e., the gluon develops a mass gap, see Alkofer:2000wg , Sternbeck:2005tk , Aguilar:2006gr , Braun:2007bx , Fischer:2008uz , Binosi:2009qm , Cyrol:2017ewj , Huber:2018ned , Falcao:2020vyr , Eichmann:2021zuv , Ferreira:2023fva and references therein.
@@
Hadrons appear as poles in QCD’s correlation functions, like the qqqqqq correlation function in the second row of Fig. 4.3. An efficient way to extract the mass of a hadron, like e.g. a nucleon resonance, is again the Bethe-Salpeter equation (BSE) discussed in Sec. 4.2, but now formulated in terms of quarks and gluons. One writes down a scattering equation $\mathbf{T}=\mathbf{K}+\mathbf{K}\,\mathbf{G}_{0}\,\mathbf{T}$, where $\mathbf{T}$ is now the connected and amputated part of the qqqqqq correlation function (i.e., without external quark propagators). $\mathbf{G}_{0}$ is the product of three dressed quark propagators which depend on the quark mass function in Fig. 4.12, so the quarks are also no longer onshell particles with a simple mass pole and a three-dimensional reduction is no longer possible. Because $\mathbf{T}$ is amputated, its residue at a given hadron pole is the hadron’s Bethe-Salpeter amplitude $\mathbf{\Gamma}$ defined through $\Psi=\mathbf{G}_{0}\,\mathbf{\Gamma}$, where $\Psi$ is the Bethe-Salpeter wave function (BSWF) from Eq. (4.8). Comparing the residues on both sides of the scattering equation then yields the homogeneous BSE at the pole,
@@
The wave function has three Dirac indices for the quarks and one for the baryon, and depending on the spin $J$ there may also be further Lorentz indices. It depends on three independent momenta, e.g. two relative Jacobi momenta $p$ and $q$ and the total momentum $P$. The wave function can be expanded in a Lorentz-covariant tensor basis, which yields 64 linearly independent tensors $\tau_{i}(p,q,P)$ for $J=1/2$ states and 128 for $J=3/2$ states Eichmann:2009qa , Sanchis-Alepuz:2011egq . From three independent momenta one can form six Lorentz invariants, one of them being onshell ($P^{2}=m^{2}$), so the dressing functions $f_{i}$ depend on five momentum variables.
@@
What are the new effects compared to the quark-model wave functions discussed in Sec. 3.2? One can arrange the $\tau_{i}$ in Eq. (4.35) into eigenstates of the three-quark spin $S$ and orbital angular momentum $L$ in the baryon’s rest frame, which results in tensors with $L=0$ ($s$ waves), $L=1$ ($p$ waves), $L=2$ ($d$ waves), and so on Eichmann:2016yit . These components mix together in the BSWF to produce a baryon with definite $J^{P}$. For example, the nucleon flavor wave function is a doublet $\mathcal{D}_{f}$, so one can cast Eq. (4.35) in permutation-group doublets $\mathcal{D}_{i}$ to obtain a fully symmetric wave function $\Psi=\sum_{i=1}^{64}\mathcal{D}_{i}\cdot\mathcal{D}_{f}$ (modulo color). One of these components is $\psi_{1}=(\mathcal{S}_{o}\,\mathcal{D}_{s})\cdot\mathcal{D}_{f}$ from Fig. 3.1, which we earlier identified with ‘the’ nucleon wave function. However, there are other $L=0$ contributions which are similarly important, and relativistic $p$ waves with $L=1$ contribute another substantial fraction of about 33% to the nucleon’s normalization Eichmann:2011vu , as we will discuss below in connection with Fig. 4.14. Relativity is therefore important for light baryons, and relativistic effects even persist up to bottom baryons Qin:2018dqp .
@@
The homogeneous BSE (4.34) is equally applicable for bound states and resonances, just like Eq. (4.8) holds in both cases. To extract a resonance mass, however, one must analytically continue the eigenvalues to unphysical Riemann sheets, which can be technically challenging Eichmann:2019dts , Santowsky:2020pwd . The situation is also more involved than with EFTs, where the particles in the loops are hadrons and so the multihadron cuts are automatic. In Fig. 4.13, the particles in the loops are quarks and gluons, so the BSE must simultaneously create hadron poles dynamically from the quark level but also the multihadron cuts, which create the Riemann sheet structure to allow for resonances, all while avoiding decay channels for quarks and gluons as required by confinement. Such a mechanism has been demonstrated in Ref. Eichmann:2015cra : A four-quark BSE implementing gluon exchanges dynamically creates intermediate meson poles, and these automatically also generate meson-meson cuts which turn a four-quark ($qq\bar{q}\bar{q}$) state into a resonance.
@@
Figure 4.14: (color online) Top: Light baryon spectrum from the three-quark and quark-diquark calculations for different $J^{P}$ quantum numbers Eichmann:2016hgl , Eichmann:2016nsu , compared to the RPP masses (real parts of pole positions). Bottom: Orbital angular momentum contributions to each baryon from the quark-diquark calculation. The bars sum up to 100%, and the light-colored boxes drawn around them are the quark-model values for $L$ from Fig. 3.2.
@@
What is the kernel $\mathbf{K}$? This is where the information from QCD’s $n$-point function enters. The leading term in the irreducible three-body kernel in Fig. 4.13, which is the three-gluon vertex connecting the three quarks, vanishes by the color traces. This hints at quark-quark correlations as the dominant forces in baryons. For the two-body kernels, most studies so far have used the rainbow-ladder truncation, where the BSE kernel turns into an effective gluon exchange Maris:1997tm , Maris:1999nt , Qin:2011dd . This preserves chiral symmetry and its dynamical breaking pattern: The quark DSE dynamically generates quark masses like in Fig. 4.12, and the pseudoscalar mesons calculated from their BSEs are the massless Goldstone bosons in the chiral limit. In the light-meson sector, rainbow-ladder works well for pseudoscalar and vector mesons, whereas scalar and axialvector mesons are too strongly bound. Going beyond rainbow-ladder by implementing higher $n$-point functions is computationally demanding, but efforts have been made and improved the meson spectrum Chang:2011ei , Williams:2015cvx , Eichmann:2016yit , Huber:2020ngt .
@@
Concerning the baryon spectrum, one strategy has been to solve the three-body equation in rainbow-ladder without further approximations Eichmann:2009qa , Eichmann:2011vu , Sanchis-Alepuz:2011egq , Sanchis-Alepuz:2014sca , Qin:2018dqp , Qin:2019hgk , Yao:2024ixu . The other is to simplify the three-body equation to a quark-diquark BSE shown in the right panel of Fig. 4.13, where gluons no longer appear and the interaction becomes a ping-pong quark exchange, with quarks forming (non-pointlike) diquarks in the process Oettel:1998bk , Eichmann:2008ef , Eichmann:2016hgl , Eichmann:2016yit , Chen:2017pse , Chen:2019fzn , Barabanov:2020jvn , Liu:2022ndb . Fig. 4.14 shows the resulting light baryon spectrum Eichmann:2016hgl , which is very similar in both cases. In particular, the nucleon mass comes out at $940$ MeV in both approaches without any tuning. Diquark clustering is a good way to think about baryons!
@@
An interesting observation in Fig. 4.14 is the correct level ordering between the Roper resonance $N(1440)\nicefrac{{1}}{{2}}^{+}$ and the nucleon’s parity partner $N(1535)\nicefrac{{1}}{{2}}^{-}$. This has not so much to do with the Roper resonance but rather with the $N(1535)$, which in contrast to the Roper is very sensitive to the higher-lying (‘ugly’) pseudoscalar and vector diquarks. In the rainbow-ladder calculation the $N(1535)$ mass comes out too low, along with many other states in the spectrum. If the strength of these ‘ugly’ diquarks is reduced, the $N(1535)$ mass moves up, overtakes the Roper resonance and ends up in the proximity of the mass of the $N(1650)$ (which is also not sensitive to the higher-lying diquarks). Corresponding plots can be found in Ref. Barabanov:2020jvn . In the three-body calculation, where diquarks do not appear explicitly, this means that the $N(1535)$ is very sensitive to effects beyond rainbow-ladder and, therefore, gluon exchange alone is not sufficient to describe the $N(1535)$ well. This is analogous to calculations in the meson sector, where the scalar and axialvector mesons receive large repulsive corrections beyond rainbow-ladder Chang:2011ei , Williams:2015cvx . Note also that the observed overall agreement with experiment in Fig. 4.14 is not in conflict with the importance of meson-baryon interactions, which are obviously necessary to generate a resonance mechanism. Unlike for higher multiquark states, such a mechanism is not automatic in the three-quark system and would enter through more advanced truncations with quark and gluon correlation functions that produce internal meson and baryon poles. The ratio $M/f_{\pi}$ (which is essentially what is shown in Fig. 4.14) would also remain unchanged if meson-baryon interactions affected the masses and $f_{\pi}$ proportionally. In any case, as expected from our discussion in Sec. 4.1, these results imply that all nucleon excitations below 2 GeV have a three-quark core.
@@
It is also interesting to study the internal composition of these baryon excitations Eichmann:2016hgl , Eichmann:2016nsu , Chen:2017pse , Liu:2022nku , Liu:2022ndb . The bars in the lower panel of Fig. 4.14 show the orbital angular momentum contributions to each baryon’s normalization. These come from the various tensors contributing to the wave functions discussed in connection with Eq. (4.35), i.e., each baryon is a mixture of different $L$ components in the rest frame. For comparison, the light-colored boxes show the quark-model values for $L$ in the harmonic oscillator assignment according to Fig. 3.2, where each baryon carries a specific value of $L$. For the majority of states one can see clear traces of the nonrelativistic quark model, e.g., the nucleon states with $J^{P}=1/2^{-}$ are dominated by $L=1$ and those with $J^{P}=3/2^{+}$ by $L=2$. But there are also clear deviations: The nucleon and $\Delta(1232)$ ground states are dominated by $L=0$ but have substantial (relativistic) $L=1$ contributions, and for several states the nonrelativistic assignment does not work at all.
@@
As will be discussed in Sec. 5, further information on the internal structure of nucleon resonances can be extracted from nucleon-to-resonance transition form factors. With functional methods, these have been studied in a variety of different approximations ranging from contact interactions Segovia:2013rca , Raya:2021pyr , Yin:2021uom , Albino:2025fcp , Albino:2025bnr , the quark-diquark approach either with model input Segovia:2015hra , Segovia:2016zyc , Chen:2018nsg , Lu:2019bjs , Chen:2023zhh or calculated rainbow-ladder ingredients Eichmann:2011vu , or in the three-quark approach Sanchis-Alepuz:2017mir . Like for the spectrum, the three-quark and quark-diquark calculations show good overall agreement, see the reviews Eichmann:2016yit , Barabanov:2020jvn . A typical common feature is missing strength at low $Q^{2}$, which in some cases can be clearly attributed to missing meson-cloud effects, e.g. for the nucleon elastic electromagnetic form factors: In that case, the meson effects are enhanced in the isovector combination but cancel in the isoscalar one, and the isoscalar anomalous magnetic moment from the three-body equation is in agreement with the experimental result $\kappa_{s}=-0.12$ Eichmann:2011vu . In other cases, discrepancies that are sometimes attributed to the meson cloud can be explained by the relativistic $L=1$ components in the baryon wave functions, like for the $R_{EM}$ ratio in the $N\to\Delta\gamma$ transition Eichmann:2011aa . For the Roper resonance, the transition form factor $F_{2}$ shows the zero crossing expected from a radial excitation, which points towards an underlying three-quark composition Segovia:2015hra . Especially important in the timelike region ($Q^{2}<0$) are the vector-meson poles, which are part of the dressed quark-photon vertex and automatically generated from the quark-gluon interactions if the vertex is calculated self-consistently Maris:1999bh . More advanced truncations that are currently being explored also capture the $\rho\to\pi\pi$ resonance mechanism which turns the $\rho$ meson into a resonance Miramontes:2019mco , Miramontes:2021xgn .
@@
Similar plots as Fig. 4.14 can be drawn for hyperons Fischer:2017cte , Eichmann:2018adq and heavy baryons Qin:2018dqp , Qin:2019hgk , Torcato:2023ijg . Especially interesting are charmed baryons such as $\Sigma_{c}$ and $\Lambda_{c}$, which consist of two light quarks and one charm quark. Denoting up and down by $n$, it is often assumed that charmed baryons form ‘planetary’ systems made of a light $(nn)$ diquark that ‘orbits’ around the heavier charm quark. However, the Faddeev calculations do not support this picture Yin:2019bxe , Torcato:2023ijg . Rather, the $(nc)n$ configurations are dominant and contribute about $60\%$ to the norm of the state, while the $(nn)c$ configurations contribute only $40\%$. Therefore, a charmed baryon actually spends most of its time in a $(nc)n$ configuration.
@@
Figure 4.15: (color online) Left: Comparison of baryon octet and decuplet masses from different lattice calculations PACS-CS:2008bkb , BMW:2008jgk , Bietenholz:2011qq , Alexandrou:2014sha , Alexandrou:2017xwd . Plot from Ref. Alexandrou:2017xwd . Right: Lattice finite-volume energy levels for $N^{*}$ and $\Delta^{*}$ excitations at a pion mass of 396 MeV Edwards:2011jj .